HUBBLE SPACE TELESCOPE DAILY REPORT #2304 PERIOD COVERED: 0000Z (UTC) 02/08/99 - 0000Z (UTC) 02/09/99 Daily Status Report as of 040/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 6967 (Blue Continuum Snapshots of 3CR Radio - Continued Galaxies) The WF/PC-2 was used to make observations of the radio galaxies 3C220.2 and 3C165.0. Radio galaxies are an important class of extragalactic objects: they represent one of the most energetic astrophysical phenomena and they may be used as probes of their environments. They are unique probes of the early Universe. The observations were completed as planned, and no problems were reported. 1.3 Completed Two Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 6353 (A Snapshot Emission Line Imaging Survey of Very Low Excitation Planetary Nebulas) The WF/PC-2 was used to make snapshot observations as part to an emission line survey of the very low excitation planetary nebula PKO38+12D1 (Class PNe). This target was selected based on the O III/H alpha line flux ratio. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Three Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The observations were completed as planned, and no anomalies were noted. 1.6 Completed STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.8 Completed STIS/CCD/MA1 7305 (The Physical Characteristics of Small-Scale ISM Structure) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations of the star HD32039 in order to measure the physical characteristics of small-scale ISM structure. A wavelength calibration observation was also taken. The observations were completed as planned, and no anomalies were reported. 1.9 Completed STIS/MA2 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.10 Completed WF/PC-2 7387 (T-Tauri Star Snapshot Survey) The WF/PC-2 was used to make observations of the stars CI-Tau and DK-Tau as part of a T-Tauri class star survey. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/MA2 7657 (STIS Spectroscopic and Imaging Sensitivity, MAMA) The Space Telescope Imaging Spectrograph (MA2) was used to make observations of GD71 in order to study the spectroscopic and imaging sensitivity. Sensitivity measurements are done for all primary tilts of the gratings and the prism used with the MAMAs, except for the G140L and G230L gratings. Imaging sensitivity measurements are done with all apertures, including a neutral density filter. A wavelength and a dark calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some high galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 7266 (High-Resolution IR and Optical Imaging of the Fields Around a Large Sample of Z > 4.5 Quasars) The WF/PC-2 was used to make observations of the quasar GB1428+421 in order to make high-resolution infrared and optical imaging of the fields around the quasars. This object appears to be very primitive and may represent early stages in the galaxy formation process. The observations were completed as planned, and no anomalies were reported. 1.14 Completed STIS/CCD 8394 (HST Service Observations of GRB 990123) The Space Telescope Imaging Spectrograph (CCD) was used to make HST service observations of GRB990123. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared once. A TTR was written for day inversions occurring on day 039. The Gyro-3 drift rate bias trend data has shown some improvement over the weekend-and-Monday period. The value achieved has been approximately 43 arcseconds/hour per day whereas the desired value of 10.8 arcseconds/hour per day is required before Gyro-3 can be re-inserted into the control loop. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH